Find the volume of the solid bounded by graphs of the equations. y + 2x - 4 = 0, y = 0, x=0
Answer #1

Ain - Find the volume of solid bounded by graph y+90-4=0 of equat ,9=0, x=0 solutions The eam y +2x-4 = 0, y=0, x=0 > · y = 4ştxf x = [6t8+ o] x š =1 LECAT] * 33/), I † -^ Sop_fat), San (bios Pop zxus = 1.08 THE FD=relem Przy : 036-KG 6 . toh tropo b